OpenStudy (vampirediaries):

Can someone help me solve these 4 problems?

OpenStudy (anonymous):

What 4???

OpenStudy (vampirediaries):

hold on please

OpenStudy (vampirediaries):

1. What are the values of a and b? 2. a) How far is the spot on the beach from the parking lot? b) How far will he have to walk from the parking lot to get to the refreshment stand?

OpenStudy (anonymous):

@Mertsj Lol!!!

OpenStudy (mertsj):

\[\frac{8}{6}=\frac{6}{a}\]

OpenStudy (mertsj):

\[\frac{18}{x}=\frac{x}{32}\]

OpenStudy (vampirediaries):

I did that and I got a= 9 over 2

OpenStudy (mertsj):

yes a=9/2

OpenStudy (vampirediaries):

How would the second equation be set up?

OpenStudy (mertsj):

Did you even read what I posted?

OpenStudy (vampirediaries):

yes.

OpenStudy (mertsj):

Then solve for x

OpenStudy (mertsj):

What did you get for x?